• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name W. Deconinck

    Log entry time 16:21:06 on October 22, 2009

    Entry number 296123

    keyword=Plan for the swing and owl shift

    The target switch over from loop 3 to loop 2 will continue until about
    6pm. The target will not be ready for cool down yet (nor tomorrow
    morning), so it will be cooled down during the Friday day shift, and cold
    by the evening.

    At 6pm (or as soon as everyone is out of the hall), we will go back to
    beam permit and we would like to follow this schedule:

    - ask MCC to do a harp scan at 2 uA
    - test whether the fast feedback (FFB) is now working
    - change to empty target and do a 100 uA test with FFB
    - change to solid target and check the raster at 5 uA with spot++ (on the
    left arm, right arm will not be ramped up yet)

    - do a Moller measurement (Sasha is notified that this is coming up)
    (the Moller measurement has precedence over FFB studies)
    - check (but not change tune) the Compton rates, while taking the Moller
    measurement (Alex should be paged when beam comes back)

    - tune the beam for Compton measurements (MCC has a procedure to tune the
    injector in this case)

    owl shift:
    - tests of oversampling on the solid target (by Bob Michaels, on shift)
    - calibration of the bcm/bpm cavities (by Zafar Ahmed, on shift)